Evaluate the expression under the given conditions:
Cos(x-y); cosx=3/5 x in quadrant 4, tan y= - square root of 3, y in quadrant 2
Side note: I’ve drawn the triangles in the appropriate quadrants but I’m not sure where to go next

Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13209)



If you have drawn the triangles correctly, then you should have
cos(x) = 3/5; sin(x) = -4/5
sin(y) = sqrt(3)/2; cos(y) = -1/2

Now plug those numbers into the formula for the cosine of the difference of the two angles:
cos(x-y) = cos(x)*cos(y)+sin(x)*sin(y)

... and evaluate.

Answer by ikleyn(52879)

.
You can find many similar problems solved in the lessons
    - Calculating trigonometric functions of angles
    - Advanced problems on calculating trigonometric functions of angles
    - Evaluating trigonometric expressions
    - OVERVIEW of lessons on calculating trig functions and solving trig equations
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Trigonometry: Solved problems".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.